Q.3A population consists of five numbers 1, 3, 4, 5, and 7. (i) Enumerate all possible samples of size two which can be drawn from the population without replacement. (ii) Calculate the mean and variance of the population. (iii) Show that the mean of the sampling distribution of the sample means is equal to the population mean. (iv) Calculate the standard error of the mean.


1
Expert's answer
2022-05-08T14:00:20-0400

We have population values 1,3,4,5 and 7 population size N=5 and sample size n=2.

(i) The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_2=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,3 & 4/2 \\
 \hdashline
 2 & 1,4 & 5/2 \\
 \hdashline
 3 & 1,5 & 6/2 \\
 \hdashline
 4 & 1,7 & 8/2 \\
 \hdashline
 5 & 3,4 & 7/2 \\
 \hdashline
 6 & 3,5 & 8/2 \\
 \hdashline
 7 & 3,7 & 10/2 \\
 \hdashline
 8 & 4,5 & 9/2 \\
 \hdashline
 9 & 4,7 & 11/2 \\
 \hdashline
 10 & 5,7 & 12/2 \\
 \hdashline
\end{array}"



(ii) Mean of population "(\mu)" = 

"\dfrac{1+3+4+5+7}{5}=4"


Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{9+1+0+1+9}{5}=4"


"\sigma=\sqrt{\sigma^2}=\sqrt{4}=2"



(iii)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X})\\ \hline
 4/2 & 1/10 & 4/20 & 16/40\\
 \hdashline
 5/2 & 1/10 & 5/20 & 25/40\\
 \hdashline
 6/2 & 1/10 & 6/20 & 36/40\\
 \hdashline
 7/2 & 1/10 & 7/20 & 49/40\\
 \hdashline
 8/2 & 2/10 & 16/20 & 128/40\\
 \hdashline
 9/2 & 1/10 & 9/20 & 81/40\\
 \hdashline
 10/2 & 1/10 & 10/20 & 100/40\\
 \hdashline
 11/2 & 1/10 & 11/20 & 121/40\\
 \hdashline
 12/2 & 1/10 & 12/20 & 144/40\\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=4=\mu"


The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{700}{40}-(4)^2=\dfrac{3}{2}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

(iv) The standard error of the mean

"\sigma_{\bar{X}}=\sqrt{\dfrac{3}{2}}\approx1.2247"